OpenStudy (anonymous):

"Solving a Polynomial Equation" x^3+64=0 So far, I have: (x)^3+64^(3) (x+64)(x^2-64x+64^2) (x+64)(x^2-64x+4096) x=-64 (?) According to the textbook example, I'm supposed to use the Quadratic formula next, which I don't understand how.

OpenStudy (anonymous):

ok, x^3 = -64 ok?

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

I don't understand at all. I know to use the "difference" of cubes" formula, \[a ^{3}+b ^{3}=(a+b)(a ^{2}-ab+b ^{2})\] to simplify. Which I did. And then simplify. And then that makes (x+64) a factor and -64 a root (I think?). But then, you have to use the Quadratic formula. And then I got to \[x=\frac{ -(-64)\sqrt{(-64)^{2}-4(x)(4096)} }{ 2a }\]

OpenStudy (anonymous):

And then \[x= \frac{ 65\sqrt{-12288} }{ 2x }\] And then I don't know. Probably got that wrong. Hate that formula.

OpenStudy (anonymous):

OH!It isn't necessary! x = -4 :)

OpenStudy (anonymous):

(-4)^3 = -64 :)

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

@wealdooo

OpenStudy (anonymous):

Oh, I know what you mean but I did that the first time I did the assignment and the instructor marked it incorrect.

OpenStudy (anonymous):

did he say why? It's quite true! :) i'm sure :)

OpenStudy (anonymous):

Well, see, I think it's cause there's a big difference between an equation and a polynomial equation. So if it was solving an equation, it'd be correct. But they want us to solve it as a polynomial (Algebra 2, by the way).

OpenStudy (anonymous):

I'm sure!but I will ask... @DebbieG ...

OpenStudy (debbieg):

Weird, I'm not familiar with that distinction, but ooookkkk.... let's run with that. so you have: \(\Large x ^{3}+4 ^{3}=(x+4)(x ^{2}-4x+4) ^{2})=(x+4)(x^2-4x+16)\) Clearly x=-4 is a solution from the first factor. The trinomial factor only has complex solutions... you could give those too, if you are looking for all real or complex solutions; but if only solving over the reals, then x=-4 is it.

OpenStudy (debbieg):

Hmm, actually, I guess I see how THAT explains the distinction. If you just apply the odd root property, you ONLY get the real solutions, not the complex ones.

OpenStudy (anonymous):

-4 isn't the only cube root of -64

OpenStudy (anonymous):

If you do it the other way, and take all of the cube roots of 64, you will get all of the answers.

OpenStudy (debbieg):

Actually, -4 is the only cube root of -64 and 4 is the only cube root of 64.

OpenStudy (anonymous):

\[ x^3 = 64e^{\pi i}\] \[ re^{ i 3\theta} = 64 e^{\pi i} \]

OpenStudy (anonymous):

Solve this complex equation to get all of the roots of 64

OpenStudy (anonymous):

-64, sorry

OpenStudy (anonymous):

Also, remember to take the angle mod 2pi

OpenStudy (anonymous):

also, I made a typo, it should be: \[r^3e^{i3\theta} = 64e^{i\pi}\]

OpenStudy (anonymous):

It is obvious that: \[r = 4\] All we have left to do is find theta

OpenStudy (anonymous):

\[3\theta = \pi + 2\pi n : n \in Z\]

OpenStudy (anonymous):

This yields the solutions: \[\theta = \pi, \frac{\pi}{3}, \frac{-\pi}{3}\]

OpenStudy (anonymous):

All other solutions are congruent to these mod 2pi

OpenStudy (anonymous):

Therefore, our solutions are: \[4e^{i\pi}, 4e^{i\pi / 3}, 4e^{-i \pi / 3}\]

OpenStudy (anonymous):

Rewriting these, we get: \[-4, 2-2i\sqrt3 , 2+ 2i\sqrt 3\]

OpenStudy (anonymous):

Which are the solutions we would get by using the sum of cubes formula.

OpenStudy (anonymous):

I'm sorry, but that's even more confusing than my textbook. :/

OpenStudy (anonymous):

Yeah, it should be

OpenStudy (anonymous):

This is much more advanced than algebra 2, and you don't need to know this

OpenStudy (anonymous):

I was just showing, that, if you take complex roots as well, either way works

OpenStudy (anonymous):

There is, in fact, no distinction between a "polynomial equation" and a "regular equation"

OpenStudy (anonymous):

Okay, well this is what my textbook is telling me to do (attatchment) But with x^3+64=0

OpenStudy (anonymous):

That would probably be the easiest way to do it.

OpenStudy (debbieg):

That's just what I said above. :) lol

OpenStudy (anonymous):

That's my problem! I don't understand the Quadratic formula, that's why I made this post. I'm just asking how to do the quadratic part. I got to \[(x+64)(x ^{2}-64x+4096)\] And I set up the quadratic part as \[x= \frac{-(-64)\sqrt{(-64)^{2}-4(x)(4096)} }{ 2x }\] and then got \[x= \frac{ 64\sqrt{-12288} }{ 2x }\] If that's even a tad bit right, I don't understand where to go from there. I do online schooling so it's really important that I understand this.

OpenStudy (debbieg):

ahhhhh, no no no. You have done the factoring wrong. No wonder you are confused!

OpenStudy (anonymous):

You did the quadratic formula right, but you didn't factor the polynomial correctly.

OpenStudy (debbieg):

I did it above, although I see I had an exponent in the wrong place. \(\Large x ^{3}+4 ^{3}=(x+4)(x ^{2}-4x+4^2)=(x+4)(x^2-4x+16)\)

OpenStudy (debbieg):

b=4. NOT b=64.

OpenStudy (anonymous):

Oh oh oh, okay! I see now. I see what I did wrong there

OpenStudy (anonymous):

So the quadratic changes to \[x= \frac{ -(-4)\pm \sqrt{(4)^{2}-4(x)(16)} }{ 2x }\] ? \[x=\frac{ 4\pm \sqrt{80} }{ 2x }\] ?

OpenStudy (anonymous):

oops, 80x

OpenStudy (debbieg):

??? the quadratic formula expression shouldn't have an x in it. The QF is a formula that SOLVES for x.

OpenStudy (anonymous):

I have no clue then. I just know, according to the textbook explanation, to use (x^2-4x+16) part of the equation for the quadratic formula

OpenStudy (anonymous):

If a polynomial is of the form: \[ax^2 + bx + c \] its roots are given by: \[\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (debbieg):

Yes, but you are using the QF wrong. If you have \(ax^2+bx+c=0\) then you compute the solutions with the QF as: \(\Large x= \dfrac{ -b\pm \sqrt{b^{2}-4ac} }{ 2a }\)

OpenStudy (debbieg):

You put in "x" where you need to put in "a".

OpenStudy (debbieg):

You aren't as far off as you think.... don't be discouraged. Just need to fine-tune that QF part. :) You have \(\Large x^2-4x+16=0\) So a=? You have b=-4 and c=16 correct.

OpenStudy (anonymous):

so is a = 1?

OpenStudy (debbieg):

exactly :)

OpenStudy (anonymous):

so is a = 1?

OpenStudy (anonymous):

\[x=\frac{ -(-4)\pm \sqrt{(4)^{2}-4(1)(16)} }{ 2(1) }\] \[x=\frac{ 4\pm \sqrt{80} }{ 2 }\]

OpenStudy (debbieg):

under the radical.... 16 - 64 = ? Then you'll be able to simplify the radical part.

OpenStudy (debbieg):

(remember, we are expecting these to be complex solutions)

OpenStudy (anonymous):

Oops. -48 not 80

OpenStudy (debbieg):

right now you can simplify that radical....

OpenStudy (anonymous):

In the textbook, the next step uses i Example: \[\frac{ 3\pm \sqrt{-27} }{ 18 }\] \[=\frac{ 3\pm3i \sqrt{3} }{ 18 }\] \[=\frac{ 1\pm i \sqrt{3} }{ 6 }\]

OpenStudy (anonymous):

So do I do that. And if so, with 3 or 4?

OpenStudy (anonymous):

or would it be 2?

OpenStudy (anonymous):

2i

OpenStudy (debbieg):

I don't know what you mean, "with 3 or 4".... just simplify the expression you have. Start with the radical. \(\sqrt{-48}\) = ??

OpenStudy (anonymous):

got -6.93

OpenStudy (debbieg):

hm. weeeeird. It's kinda like you used a calculator or something. ;) And entered \(\sqrt{48}\), took the result, and slapped a negative sign on that. Uh.uh. no no no. :) SIMPLIFY the RADICAL.

OpenStudy (anonymous):

I don't know what a radical is.....

OpenStudy (debbieg):

Factor what's under the radical sign: \(\sqrt{-48}=\sqrt{-1\cdot 16 \cdot3}\) The radical is the square root sign.

OpenStudy (debbieg):

Now anything under the radical that is a perfect square, the square root comes outside, right? e.g. \(\sqrt{18}=\sqrt{2\cdot 9}=\sqrt{2}\sqrt{9}=3\sqrt{2}\) and \(\sqrt{-18}=\sqrt{-1\cdot 2\cdot 9}=\sqrt{-1}\sqrt{2}\sqrt{9}=3i\sqrt{2}\) Now you do that, for \(\sqrt{-48}\)

OpenStudy (anonymous):

I got \[\frac{ 2\pm i \sqrt{7} }{ 1 }\] as root

OpenStudy (debbieg):

You are not going one step at a time. That's not correct, but I need you to stay with me, one step at a time, so see where your error is. I think you are having trouble simplifying the square root. You have: \(\sqrt{-48}=\sqrt{-1\cdot 16 \cdot3}\) How does that simplify?

OpenStudy (anonymous):

I did do one step at a time. I did it by the books steps though.

OpenStudy (debbieg):

I don't know what you mean, "by the book's steps". I'm sure the book is not telling you anything different than I am; it's just that the example in the book has different numbers in it. Can you show me, step by step, how you simplified \(\sqrt{-48}\) ? Because there SHOULD be an i outside, as you have, but not a 7 inside.

OpenStudy (debbieg):

Here is what "the book did": \(\Large \sqrt{-27} =\sqrt{-1\cdot9\cdot3} =\sqrt{-1}\cdot \sqrt{9}\cdot \sqrt{3} =i\cdot 3\cdot \sqrt{3} =3i\sqrt{3} \) That's how they simplified the radical part from the quad. formula. Then simplified the whole larger expression by reducing. But first, you MUST understand how to simplify that radical part. Your PROCESS should follow the book's process exactly, but your numbers are different because what's under your radical sign is different.

OpenStudy (anonymous):

OH, @DebbieG , @wealdooo , @MrMoose , I'm confused too :D so what is the answer?we should use quadratic formula or we should say that -4 is the answer?

OpenStudy (debbieg):

If you are only seeking solutions over the real numbers, then the odd root property will suffice and x=-4 is the only solution. But there are also 2 complex solutions, the roots of the quadratic part obtained from factoring the sum of cubes: \(\Large x ^{3}+4 ^{3}=(x+4)\color{red}{(x^2-4x+16)}\) You don't get those complex solutions if you only apply the odd root property, you only get them if you factor and apply the zero factor property (or, use @mrmoose's approach, but that is beyond the scope of an Algebra 2 class; factoring and using the QF, is not).

OpenStudy (anonymous):

oh thank you!so we have 3 answers!

OpenStudy (debbieg):

over the set of complex numbers, yes. :)